# Visual Affordances: Enabling Robots to Understand Object Functionality

### v1

@@ Abstract
Human-robot interaction for assistive technologies relies on the prediction of affordances, which are the potential actions a robot can perform on objects. Predicting object affordances from visual perception is formulated differently for tasks such as grasping detection, affordance classification, affordance segmentation, and hand-object interaction synthesis. In this work, we highlight the reproducibility issue in these redefinitions, making comparative benchmarks unfair and unreliable. To address this problem, we propose a unified formulation for visual affordance prediction, provide a comprehensive and systematic review of previous works highlighting strengths and limitations of methods and datasets, and analyse what challenges reproducibility. To favour transparency, we introduce the Affordance Sheet, a document to detail the proposed solution, the datasets, and the validation. As the physical properties of an object influence the interaction with the robot, we present a generic framework that links visual affordance prediction to the physical world. Using the weight of an object as an example for this framework, we discuss how estimating object mass can affect the affordance prediction. Our approach bridges the gap between affordance perception and robot actuation, and accounts for the complete information about objects of interest and how the robot interacts with them to accomplish its task.
Index Terms:
Affordance, Semantic Segmentation, Object Detection, Mass Estimation
@@ I Introduction
Affordances are the potential actions that objects in the scene offer to an agent [^1] Person as an agent is the most intuitive reference in this definition. Nevertheless, the reference to a robot is also commonly used, especially in robotics research, and the one we consider in this work.[1]. Because of such a broad definition, the computer vision and robotics research communities cast the prediction of affordances into different formulations such as grasping detection, affordance classification, affordance segmentation, and hand-object interaction synthesis [2, 3, 4, 5]. Each of these formulations addresses a part of the affordance prediction problem. The perception of affordances enables a robot to accomplish a task, selecting which objects in the environment to interact with, what actions to perform and how to perform these actions. In human-robot collaborations, considering objects and actions that are not harmful increases the challenge of affordance prediction [6, 7].
Refer to caption Refer to caption Refer to caption
Figure 1: Examples of applications benefiting from visual affordance prediction [8]: robot alone (left image), human-robot collaboration (middle image), wearable robotics and human-to-human collaboration mediated by wearable robotics (right image).
@@
Learning to perceive object affordances from visual data is challenging due to the varying appearance of objects based on the setting (e.g. single object on a tabletop or presence of clutter), the limited size of training datasets, and the characteristics of the end-effector affecting the possible interactions the robot can perform with the object. The observation of an object is affected by environmental conditions, such as illumination, background and clutter, camera viewpoint and distance, and object material (e.g. reflective), appearance (e.g. transparency or texture), and geometry (e.g. size or shape). While estimating different physical properties of an object is difficult from an image, no less is estimating the visual affordances of that object. For example, object geometry and affordance are related: concave shapes afford the holding of a content, or sharp regions afford cutting [9, 10, 11]. An object of interest can be observed as occluded by other objects in cluttered scenes [12, 13, 14] or by a human hand during a manipulation [5, 15, 16, 17]. Assistive robotics can target a specific set of objects (e.g. household containers) whose physical properties can vary during a manipulation (emptying or filling the object), or the object appearance might change due to transparencies or deformability [18] (see Fig. 2).
Most of the datasets for visual affordance have only a few tens of thousands images due to manual annotation of the object affordances, usually limited to a pre-defined set of object categories, object instances, or affordance classes [9, 12, 19]. Small-size datasets and pre-defined classes limit the generalization capabilities of learning-based models trained on these datasets. Only a few datasets [13, 20] exceed 100,000 images, with affordances automatically annotated: off-the-shelf methods [13] or simulators [20]. Training models on dataset annotated with off-the-shelf methods, however, presents challenges such as generalization to different actions and objects, since the only action annotated was the grasping of object handles [13]. Training models on datasets annotated using simulators limits the generalisation to real images due to the presence of mixed-reality images (sim-to-real gap) [20]. Different environments (or context) can imply different affordances for the same object. For example, a screwdriver can be used to insert or remove screws in a laboratory through the graspable handle. In an environment where the object does not belong to (e.g. kitchen), the screwdriver can be used to tidy up the room and the whole surface of the object becomes graspable to be picked. The physical characteristics of a robot end-effector, such as size, degrees of freedom, and number of fingers, can influence the interaction with the object. Given the same object and task, two different two-finger end-effectors (or grippers) can interact with the object in different ways. For example, to pick-up a cup of diameter 5 cm, a gripper with a maximum opening of 2.5 cm can grasp the cup only from the rim. On the contrary, to pick-up the same cup, a gripper with 5.5 cm opening can grasp also the lateral surface of the cup body.
Figure 2: Challenges in visual affordance prediction. First row: different object and different categories. Second row: different objects belonging to the same category type (intra-category diversity). Overall, objects can be captured in different scenes, backgrounds, different illumination conditions, under different poses, different occlusions, and from different viewpoints and distances. Images, cropped for visualisation purpose, are sampled from the UMD [9], IIT-AFF [12], and CCM datasets [15].
In robotics, most of the methods [2, 21, 22, 23, 24, 25] performing grasping considered the task as functional to only object picking, while ignoring the grasping as one interaction functional to different tasks (see the above broad definition of affordance). In this work, we align with the affordance definition related to the functional interaction with an object predicted from a visual input [9, 26, 27]. Given a task the robot has to perform, we consider as a visual affordance the combination of the following three aspects (what, where, how): the potential action on the most suitable objects in the image to accomplish the task; the region where the robot will interact with the object through its end-effector; and, the most physically plausible pose of the end-effector to interact with the object. All three aspects are conditioned to the task the robot needs to accomplish and therefore only one action, one region, and one pose are possible instead of multiple candidates. Hassanin et al.’s survey [26] analyses the challenges of predicting affordances from a visual input, discusses methods for affordance categorization, detection, and segmentation, and conceptually relates affordance understanding to function understanding. However, methods are only compared in terms of their architectures without discussing the limitations of approaches and datasets. Chen et al.’s survey [28] extends the number of tasks and methods comparing the backbones and performance, discusses more in depth the relations among affordance subtasks, and describes the datasets in terms of modalities, object categories, and affordance categories. Nevertheless, the discussion of previous methods focuses on classifying and segmenting affordances without analysing methods that predict the end-effector pose enabling the interaction between the robot and the object. Ardón et al.’s survey [29] classifies previous robotic methods based on the amount of prior knowledge used to build the relationships between affordance components (target object, action, and action effects), discussing the generalisation to unseen objects or environments, and the characteristics of main robotic datasets. Although the analysis provides a roadmap on how to include the concept of affordance in robotic problems, the formulations and the connections between the robotic problems are not discussed. Despite providing a comprehensive comparison, these surveys do not discuss the inconsistencies of training setups that undermine the reproducibility and fair comparison of affordance methods. Moreover, none of the previous surveys discusses the limitations in the formulation of each tasks and provides a unifying view of visual affordance that enables a robot to interact with objects.
@@
In this article, we unify the formulation for visual affordance prediction across its various tasks that were treated separately or appear disconnected in previous works and surveys. Through the lens of this unified formulation, we show its redefinition in each task and systematically review related methods and datasets, highlighting similarities and limitations of the methods and datasets. We also analyse reproducibility issues of previous works and propose an Affordance Sheet as a tool inspired by Model Cards [30] to overcome the reproducibility challenges while facilitating transparency when designing new visual affordance methods. Furthermore, we present the first framework that relates the estimation of object physical properties with visual affordance prediction to enable a robot to interact with the object in a more context-aware manner. Specifically, we show how estimating the object mass as a physical property of a manipulated container from an image is relevant for visual affordance prediction but still challenging to achieve and consequently challenging to embed into our proposed generic framework [^2] Project webpage: https://apicis.github.io/aff-survey.
Figure 3: Illustration of the unified visual affordance prediction framework for robotics. Previous works considered each block as a stand-alone affordance task unrelated from the other blocks, as also highlighted by the coloured lines. KEY:  Affordance classification,  Affordance detection and segmentation,  Affordance grounding,  Hand-object pose estimation,  Hand-object synthesis.
The article is organised as follows. We detail our unified formulation of the visual affordance prediction problem in Section II. We use our formulation as a roadmap to critically analyse the previous methods, highlighting similarities and differences, strengths and weaknesses, inside and across tasks. We analyse datasets to train and benchmark visual affordance prediction models in Section III. We discuss the performance measures to compare models in Section IV. We show a framework to estimate physical properties of objects and we discuss how the weight of an object affects the affordance prediction in Section V. In Section VI, we highlight the issue of reproducibility in previous works, highlighting characteristics and tendencies that slow down the advancements in the field along with potential solutions. We draw the conclusions in Section VII.
II Unifying the affordance problem formulation
In this section, we define a unified formulation of visual affordance prediction as the identification of the actions that a robot can perform with its end-effector on an object of interest to accomplish a task based on the visual input capturing the observed scene. We present our overall framework representing the unified formulation and comprehending the redefinitions of affordance prediction i.e., the problems tackled by previous works. We then discuss data-driven approaches that, from an RGB image, identify the object of interest in the scene, predict the object regions the robot can interact with, and predict the pose of the robot end-effector to perform the interaction.
II-A Problem formulation
Figure 4: Proposed taxonomy of tasks related to visual affordance prediction.
@@ II-A Problem formulation
Let $x_{v}\in R^{F\times W\times H\times C}$ be the observed scene, where $F$ is the number of frames of an image sequence, $W$ is the image width, $H$ is the image height, $C$ is the number of channels ($C=3$ for an RGB input), and $v$ is the camera view index in a multi-camera setup. Let $\mathcal{T}=\{t_{m}\,|\,t_{m-1}<t_{m}<t_{m+1}\}$ be a task a robot needs to perform and represented by a sequence of steps $t_{m}$ expressed as text [^3] When the cardinality of the set $\mathcal{T}$ is 1, $t_{m}$ is the task itself.. For example, a robot can be instructed to perform a task with the following steps: “close the bottle” and “move the bottle to the shelf”, or only “close the bottle”. Let $\mathcal{E}$ be the set of available end-effectors (e.g. a bi-manual robot) that can interact with the objects, and $e\in\mathcal{E}$ encode the characteristics of the end-effector (size, number of fingers, and degrees of freedom) in a parametric model (e.g. MANO [31]). Let $\mathcal{O}$ be the set of objects relevant for the task (objects of interest). Objects can be localised using an intermediate model of object detection from the image $x_{v}$ and task $\mathcal{T}$. Each object $o\in\mathcal{O}$ can be represented as a bounding box $b\in\mathbb{R}^{4}$, indicating the position and size in $x_{v}$, an object class $\lambda$, and a confidence $c$: $o=[b,\lambda,c]$. Let $\mathcal{A}_{o}$ be the set of potential actions that the robot performs on the object and each action $a\in\mathcal{A}_{o}$ can be expressed in text form. For example, for the task “close the bottle”, the robot perceives the graspable action of the bottle cap. Let $\mathcal{S}$ be the set of image regions on the object $o$ the robot can interact with to perform the action $a$. In general, to each action and object corresponds an interaction region $S^{a}_{o}\in\mathcal{S}$ on the objects of interest. $S$ can be represented as a probability map $[0,1]^{W\times H}$ having zero values in the pixels belonging to the background and values greater than zero on the object pixels. To perform the action, the robot estimates how the end-effector interacts with the object i.e., the pose of the end-effector $P$ on the interaction region of the object, also indicating how the fingers should close. For each object and action, the pose of the end-effector can be represented as a rotation-translation matrix $P=[R|T]\in SE(3)$, where $SE(3)$ is the special Euclidean group, $R\in SO(3)$ is a $3\times 3$ rotation matrix in the special orthogonal group, and $T\in\mathbb{R}^{3}$ is the translation vector in the Euclidean space. With the pose, the end-effector can be rendered on the image plane to generate $\tilde{x}_{v}\in R^{F\times W\times H\times C}$ representing the interaction between the end-effector and the object.
@@
For a given task $\mathcal{T}$ and a visual input $x_{v}$, we define a visual affordance as a region $S$ that enables a robot with its end-effector $e$ to perform an action $a$ (and $P$) on a relevant object $o$. A visual affordance model is a function that maps the observed scene $x_{v}$, the task $\mathcal{T}$, and the end-effector $e$, into the objects of interest $o$, the potential action $a$, the regions of interaction $S$, and the pose of the end effector $P$:
@@
When limited to only an RGB (RGB-D) image, we simplify the visual input to $I\in R^{W\times H\times 3}$ ($I\in R^{W\times H\times 4}$). We refer the reader to other works on affordance prediction from an RGB-D input [32, 33, 34, 35, 36, 37, 38, 39] and from multi-view inputs (including stereo) [7]. Note that, in this article, we focus on methods for visual affordance prediction from an RGB image and on the single end-effector case.
Removing the end-effector $e$, the task $\mathcal{T}$, or both, increases the number of possible solutions, making the problem too generic and irrelevant for robotics. An object can offer multiple actions for the same region, multiple regions can support the same action, and the region might not be realistic or feasible for specific agents (e.g. a robot with a 2-finger gripper). Moreover, there are additional aspects that can influence the affordance: the physical properties of the robot end-effector, the physical properties of the object, and the presence of occlusions (e.g. a person holding the object with their hand). These aspects can be included as an additional input to the visual affordance model when the problem cannot be solved, for example due to the lack of annotated data or a change in the object mass due to a manipulation or presence of a content within the object.
II-B Overall framework
We propose a framework that integrates the redefinitions related to affordance prediction (see Fig. 3) given the task to accomplish and the RGB(D) image. Our framework decomposes the task of the robot in the following subtasks and related components:
  1. 1.
    Localises the object of interest (object localisation).
  2. 2.
    Predicts the actions for each localised object (functional classification).
  3. 3.
    Predicts the object regions that enable to perform the action (functional segmentation).
  4. 4.
    Estimates the end-effector pose on the object, given the end-effector model and previous extracted information (end-effector pose estimation).
  5. 5.
    Renders the end-effector on the RGB(D) image (end-effector synthesis).
  6. 6.
    Reaches the estimated target pose, keeping into account the desired result also from a visual point of view through the rendered end-effector pose (control process).
We use our unified framework as a reference to critically analyse the previous works through the various subtasks that compose the affordance prediction problem. Each component of the framework instantiates one or more subtasks as shown in the proposed taxonomy of Fig. 4. For example, functional segmentation groups affordance segmentation and affordance grounding, as these two subtasks have a similar problem formulation; or grasping detection can be considered as a special case of end-effector pose estimation, assuming a 2-finger end-effector.
II-C Object localisation
TABLE I: Comparison of task-driven object detection and segmentation models.
@@ II-C Object localisation
Given an image $I$ and a task $\mathcal{T}$, the model predicts a set of bounding boxes $\{b_{o}\}^{O}_{o=1}$ with $b\in\mathbb{R}^{4}$ a binary segmentation mask $\{S_{o}\}^{O}_{o=1}$, with $S\in[0,1]^{W\times H}$,
@@
The challenges of task-driven object detection task lie in fusing information about object appearance and context. Some objects in the scene are not relevant for the task and different object categories can be used for the same task, thanks to the similarities in function.
We summarize the characteristics of task-driven object detection and segmentation methods in Table I. Methods tackle this task either with a single architecture trained end-to-end [40, 49] or with the combination of different models [43, 47, 51]. To understand the context and what objects are the most relevant to accomplish a task, GGNN [40] learns the dependency between objects in the scene using Graph Neural Networks, where each graph node represents an object. For each object, the model predicts the probability of being suitable for the task. However, the assumption of a closed set of tasks and objects limits the generalization to unseen objects and unknown tasks. TOIST [49] overcomes the limitation of closed set of objects for each task adapting DETR object detector [50] through a teacher-student training paradigm. The preference between objects is learned by replacing the object noun with an indefinite pronoun in the student task and by distilling the teacher knowledge from the output tokens. The representation of verb-pronoun is learned by replacing the pronoun tokens in the student with the closest noun token in the teacher using nearest neighbour. Other methods tackle the generalization to unseen objects and tasks, integrating Vision-Language Models or Large Language Models (LLM) [51, 43]. For example, CoTDet prompts an LLM to list the objects required to accomplish a task, the rationale that makes each object useful, and the (textual) features related to the rationale. Cross-attention combines the vision tokens with the textual features [51] to predict the object bounding box in the image. The alignment between objects and task can also be obtained adapting CLIP [44]. TaskCLIP applies cross-attention between vision and text tokens before the similarity, and a score function based on self-attention selects the objects that are more suitable for the task using the similarity matrix [43].
II-D Functional classification
Functional classification, also referred to as affordance classification or affordance recognition, identifies what are the potential actions (or affordance classes) $c$ that a robot can perform on an object from an input image $I$ given a task $\mathcal{T}$,

### v2

# Visual Affordance Prediction: Survey and Reproducibility

@@ Abstract
Affordances are the potential actions an agent can perform on an object, as observed by a camera. Visual affordance prediction is formulated differently for tasks such as grasping detection, affordance classification, affordance segmentation, and hand pose estimation. This diversity in formulations leads to inconsistent definitions that prevent fair comparisons between methods. In this paper, we propose a unified formulation of visual affordance prediction by accounting for the complete information on the objects of interest and the interaction of the agent with the objects to accomplish a task. This unified formulation allows us to comprehensively and systematically review disparate visual affordance works, highlighting strengths and limitations of both methods and datasets. We also discuss reproducibility issues, such as the unavailability of methods implementation and experimental setups details, making benchmarks for visual affordance prediction unfair and unreliable. To favour transparency, we introduce the Affordance Sheet, a document that details the solution, datasets, and validation of a method, supporting future reproducibility and fairness in the community.
@@ I Introduction
Affordances are the potential actions that objects in the scene offer to an agent (i.e. a human or a robot) [1]. Because of such a broad definition, the prediction of affordances is generally cast into different formulations, such as grasping detection, affordance classification, affordance segmentation, and hand-object interaction synthesis [2, 3, 4, 5]. Each redefinition addresses a part of the affordance prediction problem. For example, affordance classification identifies what actions to perform; affordance detection and segmentation localizes which objects and what regions to interact with; and grasping detection predicts the object points to perform the interaction.
@@
Learning to perceive object affordances from visual data is challenging due to the varying appearance of objects based on the setting (e.g. single object on a tabletop or presence of clutter), the limited size of datasets, and the characteristics of the agent’s hand influencing the interaction with objects. Environmental conditions, such as illumination, background and clutter, camera viewpoint and distance, influence the target object affordance. For instance, occlusions caused by other objects in cluttered scenes [6, 7, 8] or by a human hand during a manipulation [9, 10] prevent the accurate perception of the target object’s functional regions, potentially causing unintended collisions or unsafe interactions. Moreover, different environments (or contexts) imply different affordances for the same object. For example, a screwdriver can be used to insert or remove screws in a workshop through the graspable handle. In an environment where the object does not belong to (e.g. kitchen), the whole surface of the screwdriver becomes graspable to move it elsewhere. Object properties, such as material (e.g. reflective), appearance (e.g. transparency or texture), and geometry (e.g. size or shape), also influence the observation of an object and the affordance. For example, concave shapes afford the holding of a content, and sharp regions afford cutting [11, 12, 13]. The physical characteristics of a hand (human or robotic), such as size, degrees of freedom, and number of fingers, can influence the interaction with the object. A gripper with small fingers can grasp a wine glass from the stem, whereas a gripper with large fingers can grasp the glass bowl but not the stem.
Figure 1: Visual affordance prediction in case of a knife: what actions the agent performs, where the hand interacts with the object (heat map), and how the interaction is performed (hand pose for cutting). Legend:  grasp,  slide,  cut,  pierce.
Multiple methods considered grasping as functional to object picking [2, 14, 15, 16, 17, 18]. On the contrary, we base our definition of affordance on the functional interaction with an object [11, 19, 20] (see Fig. 1), considering grasping as part of an actions sequence to accomplish a higher-level task, e.g. pouring the content of a bottle implies grasping the bottle or opening a can implies grasping the tab. Given a high-level task the agent has to perform, we consider as a visual affordance the combination of the following three aspects:
  • what: the potential action on the most suitable objects in the image to accomplish the task;
  • where: the region where the agent will interact with the object through its hand; and
  • how: the most physically plausible hand pose to interact with the object.
Conditioning the task with what, where, and how, limits the potential actions, number of regions, and agent’s hand poses to a close set relevant to complete the task. Our definition also enables the grouping and comparison of previous works, showing that an incomplete formulation of the affordance (maximum two aspects) was considered. None of the previous surveys [19, 21, 22] discussed the limitations in the formulation of each task or provided a unifying view of visual affordance that enables an agent to interact with objects. Despite providing an overview of methods and datasets, previous surveys did not discuss the inconsistencies of training setups that undermine the reproducibility and fair comparison of affordance methods (see Fig. 2).
@@
In this paper, we unify the formulation for visual affordance prediction across its various tasks that were treated separately or appear disconnected in previous works and surveys. Through the lens of this unified formulation, we show the redefinition of the formulation for each task and systematically review related methods and datasets, highlighting similarities and limitations. We analyse the reproducibility issues of previous works and design the Affordance Sheet (inspired by Model Cards [23]) [^1] Project webpage: https://apicis.github.io/aff-survey to overcome the reproducibility challenges while facilitating transparency of new visual affordance methods.
Figure 2: Comparison of topics discussed in affordance surveys: Hassanin et al. [19], Chen et al. [21], Ardón et al. [22]. We propose an affordance formulation unifying previous redefinitions, we discuss the reproducibility issues and the limitations of datasets annotation preventing fair comparison across methods.
II Problem formulation
@@
Let $x_{v}\in R^{F\times W\times H\times C}$ be the observed scene, where $F$ is the number of frames of an image sequence, $W$ is the image width, $H$ is the image height, $C$ is the number of channels ($C=3$ for an RGB input), and $v$ is the camera view index in a multi-camera setup. Let $\mathcal{T}=\{t_{m}\,|\,t_{m-1}<t_{m}<t_{m+1}\}^{M}_{m=1}$ be a task the agent needs to perform and represented by a sequence of steps $t_{m}$ expressed as text. For example, a task could include the following steps: “close the bottle” and “move the bottle to the shelf”, or only “close the bottle”. Let $\mathcal{E}$ be the set of hands (human or robotic) that can interact with the objects, and $e\in\mathcal{E}$ encode the characteristics of the agent’s hand (size, number of fingers, and degrees of freedom) in a parametric model (e.g. MANO [24]). Let $\mathcal{O}$ be the set of objects relevant for the task (objects of interest). Objects can be localised using an intermediate model of object detection from the image $x_{v}$ and task $\mathcal{T}$. Each object $o\in\mathcal{O}$ can be represented as a bounding box $b\in\mathbb{R}^{4}$, indicating the position and size in $x_{v}$, an object class $\lambda$, and a confidence $c$: $o=[b,\lambda,c]$. Let $\mathcal{A}_{o}$ be the set of potential actions that the agent performs on the object and each action $a\in\mathcal{A}_{o}$ can be expressed in text form. For example, for the task “close the bottle”, the bottle cap affords the graspable action. Let $\mathcal{S}$ be the set of image regions on the object $o$ the agent can interact with to perform the action $a$. In general, to each action and object corresponds an interaction region $S_{o,a}\in\mathcal{S}$ on the objects of interest. $S$ can be represented as a probability map $0,1]^{W\times H}$ having zero values in the pixels belonging to the background and values greater than zero on the object pixels. To perform the action, the agent estimates how its hand interacts with the object, i.e. the pose of the hand $P$ on the interaction region of the object, also indicating how the fingers should close. For each object and action, the pose of the hand can be represented as a rotation-translation matrix $P=[R|T]\in SE(3)$, where $SE(3)$ is the special Euclidean group, $R\in SO(3)$ is a $3\times 3$ rotation matrix in the special orthogonal group, and $T\in\mathbb{R}^{3}$ is the translation vector in the Euclidean space. With the pose, the hand can be rendered on the image plane to visualise the interaction with the object ($\tilde{x}_{v}\in R^{F\times W\times H\times C}$).
@@
For a given task $\mathcal{T}$ and a visual input $x_{v}$, we define a visual affordance as a region $S$ that enables an agent with its hand $e$ to perform an action $a$ through a pose $P$ on a relevant object $o$. A visual affordance model is a function that maps the observed scene $x_{v}$, the task $\mathcal{T}$, and the hand $e$, into the objects of interest $o$, the potential action $a$, the regions of interaction $S$, and the pose of the end effector $P$:
@@
In this paper, we focus on methods for visual affordance prediction from an RGB image (visual input is $x_{v}=I;I\in R^{W\times H\times 3}$) and on the single hand case; we refer the reader to other works on affordance prediction from an RGB-D input [25, 26, 27, 28, 29] and from multi-view inputs (including stereo) [30]. Removing the hand $e$, the task $\mathcal{T}$, or both, increases the number of possible solutions, making the problem too generic. An object can offer multiple actions for the same region, multiple regions can support the same action, and the region might not be realistic or feasible for specific agents (e.g. a robot with a 2-finger gripper).
III Related works
TABLE I: Comparison of methods for object localisation.
@@ III Related works
Our formulation integrates the redefinitions related to affordance prediction given the task to accomplish and the RGB image. We decompose visual affordance prediction in the following subtasks and related components:
  1. 1.
    Localise the object of interest (object localisation).
  2. 2.
    Predict the actions for each localised object (functional classification).
  3. 3.
    Predict the object regions that enable to perform the action (functional segmentation).
  4. 4.
    Estimate the agent’s hand pose on the object, given the hand model and previous extracted information (hand pose estimation).
  5. 5.
    Render the hand on the RGB image (hand synthesis).
Each component of our formulation instantiates one or more subtasks. For example, functional segmentation groups affordance segmentation and affordance grounding, as these subtasks have a similar problem formulation; grasping detection can be considered as a special case of hand pose estimation. The components of our formulation provide the information about the target pose considering the desired result also from a visual point of view (rendered hand)
III-A Object localisation
Given an image $I$ and a task $\mathcal{T}$, the model predicts a set of bounding boxes $\{b_{o}\}^{O}_{o=1}$ with $b\in\mathbb{R}^{4}$, and a binary segmentation mask $\{S_{o}\}^{O}_{o=1}$, with $S\in[0,1]^{W\times H}$,
@@ III-A Object localisation
The challenges of object localisation lie in fusing object appearance with context information, as some objects in the scene are not relevant for the task, while other objects are different but have similar functionality.
Methods either use a single architecture trained end-to-end [31, 40] or a combination of different models [34, 38, 42] (see Table I). GGNN [31] predicts the probability of each detected object being suitable for the task using Graph Neural Networks, where each node represents an object. However, the assumption of a closed set of tasks and objects limits the generalization to unseen objects and unknown tasks. TOIST [40] overcomes the limitation of closed set of objects for each task through teacher-student training: the method replaces the object name in the student task sentence with an indefinite pronoun, replaces the pronoun token with the closest teacher token (nearest neighbour), and distils the teacher output. Other methods tackle the generalization to unseen objects and tasks, integrating Vision-Language Models (VLMs) or Large Language Models (LLMs) [34, 42]. For example, CoTDet prompts an LLM to list the objects required to accomplish a task, the rationale that makes each object useful, and the object (textual) features. Cross-attention combines vision and textual tokens [42] to predict the object bounding box. Alternatively, the cross-attention can combine vision and text tokens before the CLIP [35] alignment, as in TaskCLIP [34], and a score function based on self-attention selects the objects that are more suitable for the task using the similarity matrix.
III-B Functional classification
Functional classification, also referred to as affordance classification or affordance recognition, identifies what are the potential actions (or affordance classes) $c$ that an agent can perform on an object from an input image $I$ given a task $\mathcal{T}$,